An asset was purchased for $600,000, has a salvage value of $10,000, and has a useful life of 7 years. If the asset is placed in service on September 1 {the company’s fiscal year runs from October 1 to September 30}, what is its sum-of-years depreciation charge for the first year of service? What is the depreciation charge for year two of its useful life using MACRS depreciation and given it is a GDS 3-year property class.
Answer #1

In sum of digits method,we have to calculate the following

Cost= 600000

Salvage value=10000

Life = 7 years

First year =12 months.(2015-2016)

Step(1) Calculate depreciable amount

Cost of the acquisition of asset-salvage value=depreciable amount

= 600000-10000=590000

Step (2) Calculate sum of useful life

1+2+3+4+5+6+7=28

Step (3) Calculate depreciation factors

Year 1=7/28 (since you asked for first year of service only)

Step(4) Calculate depreciation for each year(year 1)

590000*7/28=147500

(2) MACRS depreciation model

Cost of the asset= 600000

Life of the asset= 3year property class

It is given that the property is a GDS 3 year class one and that we need to find its depreciation charge on year 2 of its useful life.Also since the asset is a 3 year property class asset we use 200% DB.Here we are also using half year convention (since not specified otherwise)

So depreciation for year 2 =600000*44.45/100=266700. The amount 44.45 is obtained from MACRS depreciation table.

Since it is for year 2 we take 44.45.

The 266700 used is divided by 1/2 to arrive at the right depreciation because the table includes half a year for every year 1 but since it does not know when you are going to sell your asset or dispose your asset.So you have to separately calculate half year for those years between first year.This gives us 133350.